A tool and die company is considering the purchase of a drill press with fuzzy-logic software to improve accuracy and reduce tool wear. The company has the opportunity to buy a slightly used machine for $15,000 or new one for $21,000. Because the new machine is a more sophisticated model, its operating cost is expected to be $7000 per year, while the used machine is expected to require $8200 per year. Each machine is expected to have 25 years life with 5% salvage value. Tabulate the incremental cash flow
Answer #1

Salvage value, Used machine ($) = 15,000 x 5% = 750

Salvage value, New machine ($) = 21,000 x 5% = 1,050

For both machines, in year 25 (terminal year), Operating cost will be lower by the amount of its salvage value.

Incremental cash flow (New - Old) is as follows.

Year First Cost: Old ($) First Cost: New ($) Incremental First Cost ($) Operating Cost: Old ($) Operating Cost: New ($) Incremental Operating Cost ($) Incremental Total Cost ($)
(1) (2) (3)=(2)-(1) (4) (5) (6)=(5)-(4) =(3)+(6)
0 15,000 21,000 6,000 0 6,000
1 8,200 7,000 -1,200 -1,200
2 8,200 7,000 -1,200 -1,200
3 8,200 7,000 -1,200 -1,200
4 8,200 7,000 -1,200 -1,200
5 8,200 7,000 -1,200 -1,200
6 8,200 7,000 -1,200 -1,200
7 8,200 7,000 -1,200 -1,200
8 8,200 7,000 -1,200 -1,200
9 8,200 7,000 -1,200 -1,200
10 8,200 7,000 -1,200 -1,200
11 8,200 7,000 -1,200 -1,200
12 8,200 7,000 -1,200 -1,200
13 8,200 7,000 -1,200 -1,200
14 8,200 7,000 -1,200 -1,200
15 8,200 7,000 -1,200 -1,200
16 8,200 7,000 -1,200 -1,200
17 8,200 7,000 -1,200 -1,200
18 8,200 7,000 -1,200 -1,200
19 8,200 7,000 -1,200 -1,200
20 8,200 7,000 -1,200 -1,200
21 8,200 7,000 -1,200 -1,200
22 8,200 7,000 -1,200 -1,200
23 8,200 7,000 -1,200 -1,200
24 8,200 7,000 -1,200 -1,200
25 7,150 6,250 -900 -900